OpenStudy (anonymous):

Solve for x: 3^3 * 3^x * 3^x-1= 3^12 do you just combine alike exponents??

OpenStudy (anonymous):

when multiplying the same bases, you add the exponents: 3^3 * 3^x * 3^(x-1)= 3^12 3^(3+x+x-1)=3^12 thus 3+x+x-1=12 x+x=12-3+1 2x=13-3 2x=10 x=5

OpenStudy (anonymous):

thank you, how about if the bases were differeent?